# One Look is Enough: A Novel Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation Models on High-Resolution Images

### v1

@@ 3.2 Grouped Patch Consistency Training (GPCT)
To address the boundary artifacts introduced by patch-wise refinement, we propose a simple yet effective GPCT strategy that ensures depth consistency across patch boundaries. While previous approaches [23, 24] focus only on two diagonally adjacent overlapping patches (e.g., (A, D) or (B, C) in Fig. 4) to enforce consistency constraints, our method employs four overlapping patches (A, B, C, and D in Fig. 4) simultaneously. As shown in Fig. 4-(b), we divide the training sample into overlapping patches, so that each patch overlaps with its neighbors. After refining each patch independently, we apply a depth consistency loss $\mathcal{L}_{\text{con}}$ to enforce consistency between the depth refinement results of overlapping patches:
@@
$\mathbf{D}_{\mathrm{merged}}(p)=\frac{1}{N_{o}(p)}\sum_{i\in\{A,B,C,D\}}\mathbf{D}_{\mathrm{refine}}^{i}(p),$ (3)
where $\mathbf{D}_{\mathrm{refine}}^{i}$ is the predicted depth from patch $i$ where $i=A,B,C$ and $D$ shown in Fig.4-(b), and $N_{o}(p)$ is the number of overlapping patches contributing depth estimates at location $p$. Then, we calculate the loss between merged depth $\mathbf{D}_{\mathrm{merged}}$ and GT depth $\mathbf{D}_{\mathrm{gt}}$. Since our method computes the loss using all four patches simultaneously, every backward propagation step enforces consistency at all boundaries. In contrast, the prior method [23] only applied a consistency loss on the overlapped region between two patches (e.g., (A, D) or (B, C)) at a time, leading to insufficient boundary supervision. Our approach provides a stronger supervision signal, yielding significantly improved cross-patch consistency and smooth boundary transitions during inference. By training with this GPCT strategy, our PRO model mitigates boundary artifacts without requiring additional refinement during inference, allowing efficient patch-wise depth estimation, as only a single refinement step per patch (i.e., “One Look”) is required.
@@ 3.3 Bias Free Masking (BFM)
To utilize dense GT depth when training depth refinement models for fine-grained details, we use a synthetic dataset such as the UnrealStereo4K [38]. However, the UnrealStereo4K dataset, which provides 4K resolutions, annotates the depths of transparent objects as the depths of their backgrounds (e.g., window objects in Fig. 3), which becomes a specific bias to the dataset. To maintain the benefits of dense depth labeling while avoiding overfitting to the dataset-specific biases, we propose Bias Free Masking (BFM) that uses the prior knowledge of the pretrained zero-shot MDE models. That is, the strategy of BFM is to exclude the unreliable regions corresponding to the incorrect GT depths for the transparent objects in the synthetic dataset (UnrealStereo4K) during training. Since the pretrained zero-shot MDE model $\Psi$ can be considered to have informative prior knowledge learned from large scale image-depth pairs, they can be used to identify the unreliable regions where $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$ might significantly deviate each other, indicating potential biases to the synthetic dataset.
  Methods Publications Runtime (s) Booster ETH3D Middle14 NuScnes DIS
@@
DepthAnythingV2 NIPS 2024 - 0.0307 0.993 0.0465 0.983 0.0307 0.994 0.0979 0.106 0.882 0.059
BoostingDepth CVPR 2021 12.7 0.0330 0.993 0.0552 0.974 0.0330 0.995 0.1035 0.115 0.870 0.170
PatchFusion P=16 CVPR 2024 3.4 0.0504 0.985 0.0735 0.956 0.0450 0.989 0.1124 0.141 0.831 0.206
PatchFusion P=177 CVPR 2024 36.8 0.0496 0.986 0.0723 0.957 0.0448 0.989 0.1050 0.139 0.833 0.189
PatchRefinerP=16 ECCV 2024 1.5 0.0348 0.989 0.0435 0.985 0.0292 0.995 0.0830 0.107 0.879 0.151
PatchRefinerP=177 ECCV 2024 16.2 0.0336 0.991 0.0430 0.985 0.0292 0.995 0.0805 0.106 0.881 0.141
@@
to guarantee that the final edge mask should not miss the regions that require depth refinements, otherwise they might be excluded simply from complementing $\mathbf{M}_{\mathrm{unreliable}}$. We define the edge mask as $\mathbf{M}_{\mathrm{edge}}=\mathbf{E}_{\mathrm{c}}\cap\mathbf{E}_{\mathrm{gt}}$. Finally, the BFM that denotes the reliable mask is constructed as $\mathbf{M}_{\mathrm{BFM}}=\textbf{M}_{\mathrm{edge}}\cup\sim\textbf{M}_{\mathrm{unreliable}}$ where $\sim\textbf{M}_{\mathrm{unreliable}}$ is the complement of $\textbf{M}_{\mathrm{unreliable}}$. This ensures that the unreliable regions due to large discrepancies between $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$ can be ignored, while preserving essential edge details for training. Consequently, the supervision can only be applied for the reliable regions via a masked loss:
@@ 4.1 Datasets and Metrics
We train recent SOTA depth refinement models and our PRO model on a synthetic dataset to get the advantage of dense GT depth for a fair comparison.
UnrealStereo4K. The UnrealStereo4K dataset [38] is a synthetic dataset comprising of stereo images with 4K resolution (2,160$\times$3,840) and pixel-wise GT disparity labels. As aforementioned, its some GT labels are inaccurate especially for transparent objects as shown in Fig. 3.
Booster. The Booster dataset [29] has high-resolution (3,008 $\times$ 4,112) indoor images with specular and transparent surfaces. We use the training set with GT depth for testing. To ensure diversity, we removed images captured from the same scenes, resulting in a test set of 38 images.
ETH3D. The ETH3D dataset [35] contains high-resolution indoor and outdoor images (6,048 $\times$ 4,032) with GT depth captured by LiDAR sensors. Among the dataset, 34 images were rotated 90 degrees counterclockwise, causing the bottom sides to appear on the right sides. Before testing, we rotate these images 90 degrees clockwise to ensure that the floor sides are correctly positioned at the bottom sides.
@@
DIS-5K. DIS-5K dataset [28] is a high-resolution image segmentation dataset that provides accurately annotated masks, making it suitable for evaluating edge accuracy.
Evaluation Metrics. We use the commonly adopted metrics for DE, Absolute Relative Error (AbsRel) and $\delta_{1}$. Additionally, to evaluate the edge quality, we adopt $\mathrm{D}^{3}\mathrm{R}$ metric on the Middlebury 2014 dataset follwing [26], and measure Boundary Recall (BR) metric on the DIS-5K dataset following [3]. We utilize Consistency Error (CE) [23] to evaluate the consistency of the DE results across patches.
Figure 5: Qulitative comparisons for patch-wise DE methods on ETH3D [35] and DIS-5k [28]. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the leftmost depth column represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.
  Metric PatchFusion PatchRefiner PRO (Ours)
@@ 4.2 Implementation Details
We adopt the pretrained DepthAnythingV2 (DA2) [47] as the baseline model ($\Psi$ in Fig. 4). PRO, PatchFusion [23], and PatchRefiner [24] are (re)trained based on DA2, while BoostingDepth [26] uses DA2 without additional training. The input resolution to $\Psi$ is fixed at $518\times 518$. When we use the GPCT strategy, adjacent patches are cropped with an overlap of 224 pixels, ensuring an overlap of $224\times h$ for horizontally adjacent patches and $224\times w$ for vertically adjacent patches, where $h$ and $w$ denote the height and width of the patch, respectively. In our BFM, the dilation kernel size is empirically set to (10, 20). Training is conducted on 7,592 UnrealStereo4K samples for 8 epochs with a batch size of 64 (with gradient accumulation), taking 10 hours on a single RTX 4090 GPU.
@@ 4.3 Performance Comparison
At the test time, each input image is divided into a $4\times 4$ grid of patches. The depth of each patch is refined individually, and the refined patches are then reassembled to generate the final depth map.
Zero-shot performance on high-resolution datasets. We evaluate four patch-wise DE models: three are depth refinement models including our PRO model, BoostingDepth [26], PatchRefiner [24], and the other one is a direct DE model that is PatchFusion [23]. For PatchFusion and PatchRefiner, Our evaluations are done with two conditions: (i) ‘$\mathrm{P}$=16’ setting that uses the same patch number as our method, and (ii) ‘$\mathrm{P}$=177’ setting, where additional patches are used for test-time ensembling. Note that $\mathrm{P}$ denotes the number of patches used to reassemble final depth maps during inference. As shown in Table 1, our PRO achieves SOTA performance across all depth metrics while maintaining the lowest inference time. Notably, our BFM prevents overfitting to synthetic dataset biases, leading to a 9.5% improvement in AbsRel on the Booster dataset that contains transparent objects. In terms of edge quality, our method performs worse than BoostingDepth and PatchFusion in Boundary Recall (BR). However, as shown in the depth metrics, these two methods prioritize enhancing edge sharpness rather than focusing on enhancing overall depth accuracy. In contrast, when compared to PatchRefiner that exhibits similar depth performance, our PRO achieves superior results in terms of the edge metric (BR). In terms of consistency, our PRO achieves the best result, with an 85.9% improvement, as shown in Table 2. It demonstrates the effectiveness of our GPCT strategy in maintaining consistency between individually processed patches.
@@ 4.4 Ablation Studies
Ablation on Overlap Sizes in GPCT. Table 4 shows the effect of different overlap sizes in GPCT. Each overlap value represents the shorter side of the overlapping region between adjacent patches. As shown, increasing the overlap generally reduces the CE. However, when the overlap reaches to 448 pixels, CE begins to increase. We attribute this to the fact that highly overlapped patches capture nearly identical scene contents, resulting in fewer inconsistencies between patches. So, the depth consistency loss provides less meaningful supervision, reducing its effectiveness.
@@ 5 Conclusion
In this paper, we propose the Patch Refine Once (PRO) model for depth refinement on high-resolution images. To address the depth discontinuity problem, we introduce the Grouped Patch Consistency Training (GPCT) strategy, ensuring consistency between independently processed patches. Also, we propose Bias Free Masking (BFM) to prevent the depth refinement model from overfitting to dataset-specific biases. Through these strategies, our PRO achieves superior zero-shot performance while maintaining a low inference time (12$\times$ faster than PatchRefiner with 177 patches), outperforming recent state-of-the-art methods across diverse datasets.
APPENDIX
{strip}[Uncaptioned image]

### v2

@@ 3.2 Grouped Patch Consistency Training (GPCT)
To address the boundary artifacts introduced by patch-wise refinement, we propose a simple yet effective GPCT strategy that ensures depth consistency across patch boundaries. While previous approaches [23, 24] focus only on two diagonally adjacent overlapping patches (e.g., (A, D) or (B, C) in Fig. 4-(b)) to enforce consistency constraints, our method employs four overlapping patches (A, B, C, and D in Fig. 4-(b)) simultaneously. As shown in Fig. 4-(b), we divide the training sample into overlapping patches, so that each patch overlaps with its neighbors. After refining each patch independently, we apply a depth consistency loss $\mathcal{L}_{\text{con}}$ to enforce consistency between the depth refinement results of overlapping patches:
@@
$\mathbf{D}_{\mathrm{merged}}(p)=\frac{1}{N_{\mathrm{o}}(p)}\sum_{i\in\{A,B,C,D\}}\mathbf{D}_{\mathrm{refine}}^{i}(p),$ (3)
where $\mathbf{D}_{\mathrm{refine}}^{i}$ denotes the refined depth prediction from patch $i$ (i.e., $i\in\{A,B,C,D\}$ as shown in Fig. 4-(b)), and $N_{\text{o}}(p)$ is the number of overlapping patches that contribute a depth estimate at pixel $p$. Then, we calculate the loss between merged depth $\mathbf{D}_{\mathrm{merged}}$ and GT depth $\mathbf{D}_{\mathrm{gt}}$.
Since our method computes the loss using all four patches simultaneously, every backward propagation step enforces consistency at all boundaries. In contrast, the prior method [23] only applied a consistency loss on the overlapped region between two patches (e.g., (A, D) or (B, C)) at a time, leading to insufficient boundary supervision. Our approach provides a stronger supervision signal, yielding significantly improved cross-patch consistency and smooth boundary transitions during inference. By training with this GPCT strategy, our PRO model mitigates boundary artifacts without requiring additional refinement during inference, allowing efficient patch-wise depth estimation, as only a single refinement step per patch (i.e., “One Look”) is required.
@@ 3.3 Bias Free Masking (BFM)
To utilize dense GT depth when training depth refinement models for fine-grained details, we use a synthetic dataset such as the UnrealStereo4K [38]. However, the UnrealStereo4K dataset, which provides 4K resolutions, annotates the depths of transparent objects as the depths of their backgrounds (e.g., window objects in Fig. 3), which introduces a specific bias in the dataset. To maintain the benefits of dense depth labeling while avoiding overfitting to the dataset-specific biases, we propose Bias Free Masking (BFM) that uses the prior knowledge of the pretrained zero-shot MDE models. That is, BFM aims to exclude the unreliable regions corresponding to the incorrect GT depths in transparent objects in the synthetic dataset (UnrealStereo4K) during training. Since the pretrained zero-shot MDE model $\Psi$ can be considered to have informative prior knowledge learned from large scale image-depth pairs, the model can be used to identify unreliable regions where $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$ significantly deviate from each other, indicating potential biases to the synthetic dataset.
  Methods Publications Runtime (s) Booster ETH3D Middle14 NuScenes DIS
@@
DepthAnythingV2 [47] NIPS 2024 - 0.0307 0.993 0.0465 0.983 0.0307 0.994 0.0979 0.106 0.882 0.059
BoostingDepth [26] CVPR 2021 12.7 0.0330 0.993 0.0552 0.974 0.0330 0.995 0.1035 0.115 0.870 0.170
PatchFusion P=16 [23] CVPR 2024 3.4 0.0504 0.985 0.0735 0.956 0.0450 0.989 0.1124 0.141 0.831 0.206
PatchFusion P=177 [23] CVPR 2024 36.8 0.0496 0.986 0.0723 0.957 0.0448 0.989 0.1050 0.139 0.833 0.189
PatchRefinerP=16 [24] ECCV 2024 1.5 0.0348 0.989 0.0435 0.985 0.0292 0.995 0.0830 0.107 0.879 0.151
PatchRefinerP=177 [24] ECCV 2024 16.2 0.0336 0.991 0.0430 0.985 0.0292 0.995 0.0805 0.106 0.881 0.141
@@
to guarantee that the final edge mask should not miss the regions that require depth refinements, otherwise they might be excluded simply from complementing $\mathbf{M}_{\mathrm{unreliable}}$. We define the edge mask as $\mathbf{M}_{\mathrm{edge}}=\mathbf{E}_{\mathrm{c}}\cap\mathbf{E}_{\mathrm{gt}}$. Finally, the reliable mask for BFM is constructed as $\mathbf{M}_{\mathrm{BFM}}=\textbf{M}_{\mathrm{edge}}\cup\sim\textbf{M}_{\mathrm{unreliable}}$ where $\sim\textbf{M}_{\mathrm{unreliable}}$ is the complement of $\textbf{M}_{\mathrm{unreliable}}$. This ensures that the unreliable regions caused by large discrepancies between $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$ are ignored, while essential edge details are preserved for training. Consequently, supervision is applied only to the reliable regions via a masked loss:
@@ 4.1 Datasets and Metrics
We train recent SOTA depth refinement models and our PRO model on a synthetic dataset (UnrealStereo4K) to get the advantage of dense GT depth for a fair comparison. At test time, we evaluate its zero-shot performance on four real datasets to demonstrate generalizability.
UnrealStereo4K. The UnrealStereo4K dataset [38] is a synthetic dataset comprising stereo images with 4K resolution (2,160$\times$3,840) and pixel-wise GT disparity labels. As aforementioned, its some GT labels are inaccurate especially for transparent objects, as shown in Fig. 3.
Booster. The Booster dataset [29] has high-resolution (3,008 $\times$ 4,112) indoor images with specular and transparent surfaces. We use the training set with GT depth for testing. To ensure diversity, we remove images captured from the same scenes, resulting in a test set of 38 images.
ETH3D. The ETH3D dataset [35] contains high-resolution indoor and outdoor images (6,048 $\times$ 4,032) with GT depth captured by LiDAR sensors. In the dataset, 34 images were rotated 90 degrees counterclockwise, causing the bottom of the scenes to appear on the right. Before testing, we rotate these images 90 degrees clockwise to ensure that the floor sides are correctly positioned at the bottom sides.
@@
DIS-5K. The DIS-5K dataset [28] is a high-resolution image segmentation dataset that provides accurately annotated masks, making it suitable for evaluating edge accuracy.
Evaluation Metrics. We use the commonly adopted metrics for DE, Absolute Relative Error (AbsRel) and $\delta_{1}$. Additionally, to evaluate the edge quality, we adopt $\mathrm{D}^{3}\mathrm{R}$ metric on the Middlebury 2014 dataset following [26], and measure Boundary Recall (BR) metric on the DIS-5K dataset following [3]. We utilize Consistency Error (CE) [23] to evaluate the consistency of the DE results across patches.
Figure 5: Qualitative comparisons for patch-wise DE methods on ETH3D [35] and DIS-5K [28]. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the leftmost depth column represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.
  Metric PatchFusion [23] PatchRefiner [24] PRO (Ours)
@@ 4.2 Implementation Details
We adopt the pretrained DepthAnythingV2 (DA2) [47] as the baseline model ($\Psi$ in Fig. 4). PRO, PatchFusion [23], and PatchRefiner [24] are (re)trained based on DA2, while BoostingDepth [26] uses DA2 without additional training. The input resolution to $\Psi$ is fixed at $518\times 518$. When we use the GPCT strategy, adjacent patches are cropped with an overlap of 224 pixels, ensuring an overlap of $h\times 224$ for horizontally adjacent patches and $224\times w$ for vertically adjacent patches, where $h$ and $w$ denote the height and width of the patch, respectively. In our BFM, the dilation kernel size is empirically set to (10, 20). Training is conducted on 7,592 UnrealStereo4K samples for 8 epochs with a batch size of 64 (with gradient accumulation), taking 10 hours on a single RTX 4090 GPU. Additional details are described in supplementary material.
@@ 4.3 Performance Comparison
At test time, each input image is divided into a $4\times 4$ grid of patches. The depth of each patch is refined individually, and the refined patches are then reassembled to generate the final depth map.
Zero-shot performance on high-resolution datasets. We evaluate four patch-wise DE models: three are depth refinement models including our PRO model, BoostingDepth [26], PatchRefiner [24], and the other one is a direct DE model that is PatchFusion [23]. For PatchFusion and PatchRefiner, we conduct evaluations under two settings: (i) ‘$\mathrm{P}$=16’ setting that uses the same patch number as our method, and (ii) ‘$\mathrm{P}$=177’ setting, where additional patches are used for test-time ensembling. Note that $\mathrm{P}$ denotes the number of patches used to reassemble final depth maps during inference. As shown in Table 1, our PRO achieves SOTA performance across all depth metrics while maintaining the lowest inference time. Notably, our BFM prevents overfitting to synthetic dataset biases, leading to a 9.5% improvement in AbsRel on the Booster dataset that contains transparent objects. In terms of edge quality, our method performs worse than BoostingDepth and PatchFusion in Boundary Recall (BR). However, as shown in the depth metrics, these two methods prioritize enhancing edge sharpness rather than focusing on enhancing overall depth accuracy. In contrast, when compared to PatchRefiner that exhibits similar depth performance, our PRO achieves superior results in terms of the edge metric (BR). To further demonstrate the generalizability of our method, we include additional experiments using alternative base models such as MiDaS v3.1 [2] and DepthAnythingV1 [46], as detailed in the supplementary material.
In terms of consistency, our PRO achieves the best result, with an 85.9% improvement, as shown in Table 2. It demonstrates the effectiveness of our GPCT strategy in maintaining consistency between individually processed patches.
@@ 4.4 Ablation Studies
Ablation on Overlap Sizes in GPCT. Table 4 shows the effect of different overlap sizes in GPCT. Each overlap value denotes the size (in pixels) of the square overlapping region shared between adjacent patches. As shown, increasing the overlap generally reduces the CE. However, when the overlap reaches 448 pixels, CE begins to increase. We attribute this to the fact that highly overlapped patches capture nearly identical scene contents, resulting in fewer inconsistencies between patches. So, the depth consistency loss provides less meaningful supervision, reducing its effectiveness.
@@ 5 Conclusion
In this paper, we propose the Patch Refine Once (PRO) model for depth refinement on high-resolution images. To address the depth discontinuity problem, we introduce the Grouped Patch Consistency Training (GPCT) strategy, ensuring consistency between independently processed patches. In addition, we propose Bias Free Masking (BFM) to prevent the depth refinement model from overfitting to dataset-specific biases. Through these strategies, our PRO achieves superior zero-shot performance while maintaining a low inference time (12$\times$ faster than PatchRefiner with 177 patches), outperforming recent state-of-the-art methods across diverse datasets.
Acknowledgements. This work was supported by IITP grant funded by the Korea government(MSIT) (No.RS2022-00144444, Deep Learning Based Visual Representational Learning and Rendering of Static and Dynamic Scenes).
[…]
Supplementary Material
{strip}
  Base Methods Booster ETH3D Middle14 DIS CE $\downarrow$
[…]
MiDaS v3.1[2] MiDaS v3.1 [2] 0.0538 0.973 0.0655 0.965 0.0399 0.991 0.1914 0.028 -
PatchFusion P=16 [23] 0.0613 0.964 0.0826 0.952 0.0465 0.989 0.1435 0.100 0.456
PatchRefiner P=16 [24] 0.0520 0.977 0.0594 0.971 0.0377 0.993 0.1396 0.058 0.461
PRO 0.0513 0.977 0.0578 0.972 0.0370 0.993 0.1261 0.062 0.077
DepthAnythingV1 [46] DepthAnythingV1 [46] 0.0482 0.977 0.0489 0.981 0.0315 0.995 0.1546 0.023 -
PatchFusion P=16 [23] 0.0621 0.962 0.0672 0.970 0.0410 0.993 0.1246 0.081 0.393
PatchRefiner P=16 [24] 0.0507 0.977 0.0480 0.982 0.0301 0.996 0.1240 0.043 0.409
PRO 0.0500 0.978 0.0455 0.983 0.0294 0.996 0.1193 0.052 0.073
Table 5: Generality across various base models. Quantitative comparison of depth estimation methods on Booster [29], ETH3D [35], Middlebury 2014 [34], and DIS-5K [28] datasets. Bold indicates the best performance in each metric.
  Meth BFM Booster ETH3D Middle14 DIS
[…]
PF [23] 0.0504 0.985 0.0735 0.956 0.0450 0.989 0.1124 0.206
0.0446 0.991 0.0761 0.956 0.0447 0.991 0.1149 0.198
PR [24] 0.0348 0.989 0.0435 0.985 0.0292 0.995 0.0830 0.151
0.0307 0.994 0.0432 0.985 0.0290 0.995 0.0791 0.151
Table 6: Effect of Bias Free Masking (BFM) on training of PatchFusion (PF) [23] and PatchRefiner (PR) [24].
Appendix A Extra Experiments
Generality across Base Models. Replacing the base model (BM) requires retraining. To show the generality of our method, we provide additional comparisons using MiDaS v3.1 [2] and DepthAnythingV1 (DA1) [46] as base models. All evaluations are conducted in the same way in the Section 4.3. As shown in the Table 5, the MiDaS v3.1- and DA1-based PRO consistently outperform the previous SOTA models. However, all DA1-based methods show inferior AbsRel on the Booster compared to the DA1. This is because, as shown in the Fig. 6, DA1 yields poor predictions on transparent objects, thus leading the other methods to further degraded performance.
Figure 6: Depth estimation results from different base models (MiDaS v3.1 [2], DepthAnythingV1 [46], and DepthanythingV2 [47].
Effect of Bias Free Masking (BFM). To further demonstrate the effectiveness of BFM, we apply it to both PatchFusion [23] and PatchRefiner [24] during training. The consistent performance improvements observed across most metrics validate the generalizability of our approach, as summarized in Table 6.
  P Booster ETH3D Middle14 DIS
[…]
$3\times 3$ 0.0305 0.994 0.0423 0.985 0.0288 0.995 0.0805 0.125
$4\times 4$ 0.0304 0.994 0.0422 0.985 0.0287 0.996 0.0803 0.156
$5\times 5$ 0.0305 0.994 0.0423 0.985 0.0288 0.996 0.0797 0.162
Table 7: Effect of patch numbers on the performance.
  Base CE $\downarrow$ Booster ETH3D Middle14 DIS
[…]
Diag 0.075 0.0302 0.994 0.0423 0.986 0.0288 0.996 0.0787 0.171
GPCT 0.049 0.0304 0.994 0.0422 0.985 0.0287 0.996 0.0803 0.156
Table 8: Quantitative comparisons between regularizing consistency diagonally (Diag) and GPCT.
Effect of Patch numbers on Performance. To investigate the effect of the number of patches during inference, we conduct experiments with the PRO model by partitioning the input images into 3$\times$3, 4$\times$4, and 5$\times$5 patches, and compare their performance. As we can see in the Table 7, depth metrics remain stable across patch numbers, while edge metrics improve with more patches. This allows users to adjust patch counts at inference to balance speed and detail without retraining.
Ablation on GPCT vs Diagonal consistency. We additionally provide comparison between regularizing 2-diagonal patches and our GPCT in the Table 8. While the depth metrics show marginal differences, GPCT achieves significantly better consistency in CE metric qualitatively as shown in the Fig.7.
Figure 7: Qualitative comparisons between Diagonal approach and GPCT.